OpenStudy (tumblewolf):

Will medal! Give intervals where f is concave up and the intervals where f is concave down. Identify all inflection points. F(x)= 2x^6-3x^5-10x^4 I have my F'' being 60x^4-60x^3-120x^2

OpenStudy (agent0smith):

You can factor out an x^2, then set it equal to zero to find critical values.

OpenStudy (tumblewolf):

Don't you need a denominator? To find where it doesn't exist?

OpenStudy (anonymous):

After factoring out 60x^2, you are left with\[60x^2(x^2-x-2)=0\] Then solve for x, to find concavity of f(x) using f''(x).

OpenStudy (agent0smith):

Polynomials exist everywhere. Use the same reasoning as the last question. Find critical values, make a number line. Concave up means f'' is positive, concave down means f'' is negative. Inflection points are where the f'' changes sign.

OpenStudy (tumblewolf):

-1 and 2?

OpenStudy (agent0smith):

You missed one.

OpenStudy (anonymous):

-1 and 2 are correct for setting x^2-x-2=0 60x^2=0 divide both sides by 60, x^2=0 sqrt both sides, x=0

OpenStudy (tumblewolf):

I used F''(-2) F''(-0.5) F''(1) F''(3) And got all increasing

OpenStudy (tumblewolf):

Of concave up

OpenStudy (tumblewolf):

So are there no inflection points?

OpenStudy (agent0smith):

I'm guessing you made some mistakes. https://www.google.com/search?q=2x%5E6-3x%5E5-10x%5E4&oq=2x%5E6-3x%5E5-10x%5E4&aqs=chrome..69i57&sourceid=chrome&ie=UTF-8 That's the graph of the function. It pretty clearly has some regions where it's concave down.

OpenStudy (tumblewolf):

So what did I do wrong?

OpenStudy (agent0smith):

Probably these F''(-2) F''(-0.5) F''(1) F''(3)

OpenStudy (agent0smith):

^Plugging them into f'' is correct.

OpenStudy (tumblewolf):

F'' is the 60x^2-(x^2-x-2) right?

OpenStudy (agent0smith):

Not quite... 60x^2(x^2-x-2)

OpenStudy (mathmale):

Starting with F(x)= 2x^6-3x^5-10x^4, you'll need to find both the first and the second derivatives to answer these questions completely. I'd suggest you show your work (including intermediate results).

OpenStudy (tumblewolf):

The first derivative is 12x^5-15x^4-40x^3 and the second derivative is 60x^4-60x^3-120x^2

OpenStudy (agent0smith):

If you used f''(x) = 60x^2-(x^2-x-2) then that would explain why they were all positive.

OpenStudy (tumblewolf):

I did, so what would it be?

OpenStudy (agent0smith):

It's not 60x^2-(x^2-x-2), it's 60x^2(x^2-x-2)

OpenStudy (tumblewolf):

Okay, this time I got F''(-2) - F''(-.5) - F''(1) - F''(3) +

OpenStudy (tumblewolf):

Oh I see! I was doing my math wrong in the calculator! So the inflection points are -1 and 2?

OpenStudy (agent0smith):

Yep.

OpenStudy (tumblewolf):

Oh okay thank you! I have one more that's kind of similar but I'm going to try and work it myself. Thank you for all the help and sorry for being all confusing!

OpenStudy (agent0smith):

You're welcome.

OpenStudy (tumblewolf):

@agent0smith I'm sorry! One more quick question! I can only find one point in a second derivative test and can't find another

OpenStudy (agent0smith):

Why do you need more than one point?

OpenStudy (tumblewolf):

Can you only do it with one?

OpenStudy (tumblewolf):

Do it with only one

OpenStudy (agent0smith):

Yes